A hybrid car manufacturer made a vehicle that the company claimed got 29 kilometers per liter. What would be the approximate equivalent gas consumption in miles per gallon?
Answer by ankor@dixie-net.com(22740)

A hybrid car manufacturer made a vehicle that the company claimed got 29 kilometers per liter.
What would be the approximate equivalent gas consumption in miles per gallon?
:
A couple useful conversion factors:
3.785 liters in 1 gallon
.6214 miles in 1 km
:
Find miles per liter
.6214 * 29 = 18.02 mile/liter
:
Find miles per gallon
3.785 * 18.02 = 68.2 miles/gallon
:
:
A single conversion factor to convert km/liter to mi/gal
.6214 * 3.785 = 2.352
:
2.352 * 29 = 68.2